Write the chemical formula for the complex salt us How do I solve this and solutions
Answer #1

3. as per the data

x = +1

y = +3

z = -2

formula = xyz2*3H2O   = KFe(C2O4)2*3H2O

4. molarmass = 325.0271 g/mol

5. mass of salt not given